OpenStudy (anonymous):

identify the restrictions on the variable? x^2+3x+2/x^2-4x-12

OpenStudy (anonymous):

X is restricted to those values where the denominator is not equal to 0. So, just find where x^2-4x-12 = 0. Those values of x are not allowable.

OpenStudy (anonymous):

Hint: x^2-4x-12 = 0 can be factored very easily. You do not have to resort to the quadratic formula for this problem.

OpenStudy (anonymous):

so i just solve that equation and that's my answer? solve x^2-4x-12=0?

OpenStudy (anonymous):

That all that needs to be done.

OpenStudy (anonymous):

oh okay, thank you so much.

OpenStudy (anonymous):

You are quite welcome! Medals are nice!